OpenStudy (slayer8001):

The area of a trapezoid is given by the formula A = (b 1 + b 2)h, where base b 1 is parallel to base b 2 and h is the height. Solve the formula for b 2. Show your work. the answer should end up as..... b2=2A/h-b1

OpenStudy (slayer8001):

@aloud

OpenStudy (anonymous):

A1/2=(b1+b2)h A1/2=bh1+bh2 A1/2-bh1=bh2 bh2=A1/2-bh1 its b2=(A1/2-bh1)/h

OpenStudy (anonymous):

i think its 2x - 5

OpenStudy (slayer8001):

thanks guys